a triangle has 5, 6 sides and 9 cm. if a similar triangle has a perimeter of 200 cm, what are the sides of this triangle?

Answer by checkley79(3341)

5+6+9=20 INCHES FOR THE FIRST TRIANGLE
200/20=10 THIS IA THE MULTIPLIER FOR THE THREE SIDES.
5*10=50 )
6*10=60 ) NEW SIDED.
9*10=90 )
PROOF:
50+60+90=200